# Around the Merino–Welsh conjecture: improving Jackson’s inequality

### v1

@@ 2.2. Permutation Tutte polynomial
$Y_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
@@
(b1) If $x\geq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function for each $i\in A$.
(b2) If $0\leq x\leq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function for each $i\in A$.
(b3) For $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
(b4) Finally, for $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing for each $j\in B$.
Remark 2.12.
This lemma is the reason why we generated $x_{j}$ by $x_{j}=1-y_{j}$ for $j\in B$. If we consider the function
$X^{\prime}_{i}(x_{i},\{x_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}>x_{i}.\end{array}\right.$
instead of $X_{i}(x_{i},\{y_{j}\}_{j\in B})$, this would be neither increasing, nor decreasing.
We will also need a very little extension of the above ideas, where each vertex gets its own activity: let For $i\in A$, let us introduce the random variable
$\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x^{(i)}&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$
@@
$\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y^{(j)}&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
The following lemma is just a trivial extension of the previous lemma together with Harris’ inequality.
@@
(a1) If $i\in A$ and $x^{(i)}\geq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function.
(a2) If $i\in A$ and $0\leq x^{(i)}\leq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function.
(a3) For $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
(a4) Finally, if $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing.
(b) If $x^{(i)}\geq 1$ for all $i\in A$ and $0\leq y^{(j)}\leq 1$, then
$\mathbb{E}\left[\prod_{i\in A}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]\geq\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]=\prod_{i\in A}\left(\frac{x^{(i)}}{d_{i}+1}+\frac{d_{i}}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(\frac{y^{(j)}}{d_{j}+1}+\frac{d_{j}}{d_{j}+1}\right).$
@@
Let $x\geq 1$ and $0\leq y\leq 1$. Let $H_{1}$ be a rooted graph with root vertex $v_{1}$. Let $H_{2}$ be another rooted graph with root vertex $v_{2}$. Let $H$ be obtained from $H_{1}$ and $H_{2}$ by identifying $v_{1}$ and $v_{2}$ in the union of $H_{1}$ and $H_{2}$. Let $v$ be the vertex obtained from identifying $v_{1}$ and $v_{2}$. Assume that the bipartite parts of $H$ determines the bipartite parts of $H_{1}$ and $H_{2}$, that is, if $v\in A(H)$, then $v_{1}\in A(H_{1})$ and $v_{2}\in A(H_{2})$, and if $v\in B(H)$, then $v_{1}\in B(H_{1})$ and $v_{2}\in B(H_{2})$.
@@
$x\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
@@
$\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\frac{1}{x}(\widetilde{T}_{H_{1}}(x,0)\widetilde{T}_{H_{1}}(0,x))(\widetilde{T}_{H_{2}}(x,0)\widetilde{T}_{H_{2}}(0,x))$
3. Warm-up: asymptotic computation of some permutation Tutte polynomial
In this section we motivate an important technique of this paper by computing the asymptotic value of the permutation Tutte polynomial of certain bipartite graphs.
@@ 3. Warm-up: asymptotic computation of some permutation Tutte polynomial
We generate the random permutation on $V(H)=A\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let $t=\max_{v\in A}x_{v}$ and $s=\max_{v\in B}x_{v}$. We have $ab$ choices for the vertices that take the value $t$ and $s$, let these vertices be $v_{A}$ and $v_{B}$. If $s>t$, then the permutation has weight 00 as $v_{B}$ is an active vertex. If $t>s$, then no vertex can be active in $B$. For each $w\in A$ that is not equal to $v_{A}$ two things can happen: if $x_{w}<s$, then $w$ is not active, so $X_{w}=1$, or $s<x_{w}<t$ and then $w$ is active so $X_{w}=x$. For a vertex $w\in B$ we simply need to have $x_{w}<s$ and $w$ will not be active. The integral formula then follows. ∎
@@
Clearly, at the maximum we have $t=1$ and we simply need to maximize $s^{\beta}(s+x(1-s))^{\alpha}$. This turns out to be at $s=\min\left(1,\frac{\beta x}{\beta-1}\right)$. If $\beta<\frac{x-1}{x}$, then the growth constant is
$\left(\frac{\beta x}{\beta-1}\right)^{\beta}\left(x+(1-x)\frac{\beta x}{x-1}\right)^{\alpha}=\alpha^{\alpha}(1-\alpha)^{1-\alpha}\frac{x}{(x-1)^{1-\alpha}}.$
If $\beta\geq\frac{x-1}{x}$, then the exponential growth constant is simply $1$.
@@
It turns out that for even $n$ there are matroids for which all local basis exchange graphs are isomorphic to $H_{n,n,n}$. Indeed, all local basis exchange graphs of the matroid $U^{(2)}_{\frac{3}{2}n,n}$ are isomorphic to $H_{n,n,n}$, where $U_{\frac{3}{2}n,n}$ is the uniform matroid of rank $n$ on $\frac{3}{2}n$ elements, and the matroid $M^{(2)}$ is obtained from $M$ by replacing each element of $M$ with two parallel elements. Furthermore, we have
$\widetilde{T}_{H_{n,n,n}}(2,0)\widetilde{T}_{H_{n,n,n}}(2,0)<\widetilde{T}_{H_{n,n,n}}(1,1)^{2}$
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}=\max_{s\in[0,1]}\left(s+x(1-s)\right)\left(xs+(1-x)\frac{s^{2}}{2}\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}&\text{if}\ \ x\geq\sqrt{3},\\ \frac{1}{2}(x+1)&\text{if}\ \ 1<x\leq\sqrt{3}\end{cases}$
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}=\max_{s\in[0,1]}t\left(\frac{t^{2}}{2}+\left(\frac{1}{2}-\frac{t^{2}}{2}\right)x\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}&\text{if}\ \ x\geq\frac{3}{2},\\ \frac{1}{2}&\text{if}\ \ 1<x\leq\frac{3}{2}.\end{cases}$
In particular, if $x\geq\sqrt{3}$ we have
@@
As before we generate the random permutation on $V(H)$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let
@@
Let $v_{A}$ and $v_{B}$ be the vertices, where these maximum are achieved. (Note that $t$ is only maximum on $A$ and not on $A\cup C$.) If $s<t$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
[…]
Here $s+x(t-s)$ is the contribution of a vertex $w\in A$ that is not $v_{A}$. If $x_{w}=s_{i}<s$ for some $w\in B$ that is not $v_{B}$, then the attached leaf in $C$ has contribution $s_{i}+x(1-s_{i})$. (Note that attached leaf vertex $u$ can have a value bigger than $t$ as $t$ was only a maximum among the vertices of $A$.) The last term $s+x(1-s)$ is simply the contribution of the leaf attached to $v_{B}$.
@@
[…]
Here the contribution of a vertex $w\in A$ is simply $t$ as it can never be active since $s<t$. If we have a vertex $w\in B$ that is not $v_{B}$ with $x_{w}=s_{i}<s$, then we need to distinguish two cases. If $s_{i}<t$, then $w$ is not active, and so the attached leaf has a contribution $s_{i}+x(1-s_{i})$. If $s_{i}>t$ then, in order to make $w$ inactive the attached leaf $u$ should have a value $x_{u}>s_{i}$, but then it is automatically active. The contribution of the leaf attached to $v_{B}$ is $x(1-s)$ as it has to be active to make $v_{B}$ inactive.
@@
$\lim_{n\to\infty}I_{1}^{1/n}$ $=\max_{0\leq s\leq t\leq 1}(s+x(t-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
$=\max_{0\leq s\leq 1}(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
The function $(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$ has either maximum at $\left(1-\frac{1}{\sqrt{3}}\right)\frac{x}{x-1}$ if this is less than $1$, or at $s=1$. In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}$. In the second case, its value is $\frac{1}{2}(x+1)$.
@@
$\lim_{n\to\infty}I_{2}^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left((1-x)\frac{t^{2}}{2}+xt+x(s-t)-x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
$=\max_{0\leq t\leq s\leq 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
$=\max_{0\leq s\leq 1}\left(x\left(s-\frac{s^{2}}{2}\right)+\frac{1}{2}\right)$
$=\frac{1}{2}(x+1),$
so this is always at most $\lim_{n\to\infty}I_{1}^{1/n}$.
Then we have
@@
This proves the first statement. The second statement is actually simpler. We have
[…]
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
$=\max_{0\leq t\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{1}{2}-\frac{t^{2}}{2}\right)\right)$
This expression has either maximum at $x=\left(\frac{x}{3(x-1)}\right)^{1/2}$ if $x\geq\frac{3}{2}$ or at $1$ if $x\leq\frac{3}{2}$. In the first case its value is $\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}$. In the second case, its value is $\frac{1}{2}$. This completes the sketch of the proof. ∎
@@
It is clear from the above examples that even though the permutation Tutte polynomial might be complicated still it is possible to compute the exponential growth of it for many families of bipartite graphs. Another important take-away that it is often possible to understand those permutations that contributes the most weight to the permutation Tutte polynomial, and concentrating for these permutations might give a better lower or upper bound for the permutation Tutte polynomial.

### v2

@@ 2.2. Permutation Tutte polynomial
$Z_{j}(\{x_{i}\}_{i\in A},x_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant x_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>x_{j}.\end{array}\right$.
As a consequence of Lemma 2.10 we have
(2.1) $\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
The problem with the random variables $Z_{v}(\underline{x})$ is that they are neither monotone increasing, nor monotone decreasing.
In what follows, we do a little trick. For $i\in A$ we generate $x_{i}\sim U(0,1)$ as before, but for $j\in B$ we first generate a uniform random number $y_{j}$ from $[0,1]$ and let $x_{j}=1-y_{j}$. The role of this trick will become apparent soon.
For $i\in A$, let us introduce the random variable
$X_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right$.
and for $j\in B$, let
$Y_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>1-y_{j}.\end{array}\right$.
@@
(b1) If $x\geqslant 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function for each $i\in A$.
(b2) If $0\leqslant x\leqslant 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function for each $i\in A$.
(b3) For $0\leqslant y\leqslant 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ increases monotonically for each $j\in B$.
(b4) Finally, for $y\geqslant 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ decreases monotonically for each $j\in B$.
We will also need a slight extension of the above ideas, where each vertex gets its own activity. For $i\in A$, let us introduce the random variable
$\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x^{(i)}&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right$.
@@
$\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y^{(j)}&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>1-y_{j}.\end{array}\right$.
The following lemma is just a trivial extension of the previous lemma together with Harris’s inequality.
@@
(a1) If $i\in A$ and $x^{(i)}\geqslant 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function.
(a2) If $i\in A$ and $0\leqslant x^{(i)}\leqslant 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function.
(a3) If $j\in B$ and $0\leqslant y^{(j)}\leqslant 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing.
(a4) Finally, if $j\in B$ and $y^{(j)}\geqslant 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing.
(b) If $x^{(i)}\geqslant 1$ for all $i\in A$ and $0\leqslant y^{(j)}\leqslant 1$ for all $j\in B$, then
$\mathbb{E}\left[\prod_{i\in A}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]\geqslant\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]=\prod_{i\in A}\left(\frac{x^{(i)}}{d_{i}+1}+\frac{d_{i}}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(\frac{y^{(j)}}{d_{j}+1}+\frac{d_{j}}{d_{j}+1}\right).$
@@
Let $x\geqslant 1$ and $0\leqslant y\leqslant 1$. Let $H_{1}$ and $H_{2}$ be rooted bipartite graphs with root vertices $v_{1}$ and $v_{2}$, respectively. Let $H$ be obtained from $H_{1}$ and $H_{2}$ by identifying $v_{1}$ and $v_{2}$ in the union of $H_{1}$ and $H_{2}$. Let $v$ be the vertex obtained from identifying $v_{1}$ and $v_{2}$. Assume that the bipartite parts of $H$ determines the bipartite parts of $H_{1}$ and $H_{2}$, that is, if $v\in A(H)$, then $v_{1}\in A(H_{1})$ and $v_{2}\in A(H_{2})$, and if $v\in B(H)$, then $v_{1}\in B(H_{1})$ and $v_{2}\in B(H_{2})$.
@@
$x\widetilde{T}_{H}(x,y)\geqslant\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
@@
$\widetilde{T}_{H}(x,y)\geqslant\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
@@
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\frac{1}{x}(\widetilde{T}_{H_{1}}(x,0)\widetilde{T}_{H_{1}}(0,x))(\widetilde{T}_{H_{2}}(x,0)\widetilde{T}_{H_{2}}(0,x)).$
3. Warm-up: asymptotic computation of some permutation Tutte polynomials
In this section, we motivate an important technique of this paper by computing the asymptotic values of the permutation Tutte polynomial of certain bipartite graphs.
@@ 3. Warm-up: asymptotic computation of some permutation Tutte polynomials
Let $K_{a,b}=(A,B,E)$ with $|A|=a,|B|=b$. As before we generate the random permutation on $A\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V(K_{a,b})$, and then we take the relative order of $x_{v}$’s. Let $t=\max_{v\in A}x_{v}$ and $s=\max_{v\in B}x_{v}$. Recall that from equation 2.1 we have
$\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
We will refer to $\prod_{v\in V(H)}Z_{v}(\underline{x})$ as the weight corresponding to the permutation determined by $\underline{x}$ as it only depends on the ordering of $(x_{v})_{v\in V}$. Our plan is to compress this integral using $s$ and $t$. We have $ab$ choices for the vertices that take the values $t$ and $s$, let these vertices be $v_{A}$ and $v_{B}$. If $s>t$, then the permutation has weight 0 as $v_{B}$ is an active vertex. If $t>s$, then no vertex can be active in $B$. For each $w\in A$ that is not equal to $v_{A}$ two things can happen: if $x_{w}<s$, then $w$ is not active, so $Z_{w}(\underline{x})=1$, or $s\leqslant x_{w}\leqslant t$ and then $w$ is active so $Z_{w}(\underline{x})=x$. This shows that as $x_{w}$ runs over the interval $[0,t]$ its contribution to the integral is s+x(ts)s+x(t-s). For a vertex $w\in B$ we simply need to have $x_{w}<s$ and $w$ will not be active, so we integrate $1$ on the whole interval $[0,s]$. The integral formula then follows. ∎
@@
Clearly, at the maximum we have $t=1$ and we simply need to maximize $s^{\beta}(s+x(1-s))^{\alpha}$. This turns out to be at $s=\min\left(1,\frac{\beta x}{x-1}\right)$. If $\beta<\frac{x-1}{x}$, then the growth constant is
$\left(\frac{\beta x}{x-1}\right)^{\beta}\left(x+(1-x)\frac{\beta x}{x-1}\right)^{\alpha}=\alpha^{\alpha}(1-\alpha)^{1-\alpha}\frac{x}{(x-1)^{1-\alpha}}.$
If $\beta\geqslant\frac{x-1}{x}$, then the exponential growth constant is simply $1$.
@@
It turns out that for even $n$ there are matroids for which all local basis exchange graphs are isomorphic to $H_{n,n,n}$. Indeed, let us first consider the uniform matroid $U_{n,\frac{3}{2}n}$ of rank $n$ on $\frac{3}{2}n$ elements. Then let us replace all elements of $U_{n,\frac{3}{2}n}$ with $2$ parallel elements. Let us denote the resulting matroid by $U^{(2)}_{n,\frac{3}{2}n}$. (In general, if we replace each element of a matroid $M$ with $k$ parallel elements, then we denote by $M^{(k)}$ the obtained matroid.) The resulting matroid has $3n$ elements and rank $n$. If $B$ is a basis of the matroid $U^{(2)}_{n,\frac{3}{2}n}$, then $C$ will be the set of corresponding parallel elements, and $A$ will be the remaining elements. So the local basis exchange graph will be the dual bipartite graph $H^{*}_{n,n,n}=(B,A\cup C,E)$. For the dual matroid of $U^{(2)}_{n,\frac{3}{2}n}$ the local basis exchange graph will be exactly $H_{n,n,n}=(A\cup C,B,E)$. Below we sketch an alternative proof for the fact that
$\widetilde{T}_{H_{n,n,n}}(2,0)\widetilde{T}_{H_{n,n,n}}(0,2)<\widetilde{T}_{H_{n,n,n}}(1,1)^{2}$
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}=\max_{s\in[0,1]}\left(s+x(1-s)\right)\left(xs+(1-x)\frac{s^{2}}{2}\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}&\text{if}\ \ x\geqslant\sqrt{3},\\ \frac{1}{2}(x+1)&\text{if}\ \ 1<x\leqslant\sqrt{3},\end{cases}$
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}=\max_{t\in[0,1]}t\left(\frac{t^{2}}{2}+\left(\frac{1}{2}-\frac{t^{2}}{2}\right)x\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}&\text{if}\ \ x\geqslant\frac{3}{2},\\ \frac{1}{2}&\text{if}\ \ 1<x\leqslant\frac{3}{2}.\end{cases}$
In particular, if $x\geqslant\sqrt{3}$ we have
@@
As before we generate the random permutation on
$V(H)=(A\cup C)\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let
@@
Let $v_{A}$ and $v_{B}$ be the vertices, where these maximums are achieved. (Note that $t$ is only maximum on $A$ and not on $A\cup C$.) As before we use equation 2.1:
$\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
and we aim to compress this integral using $s$ and $t$.
If $s<t$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
(3.1) $I_{1}:=n^{2}\int_{0}^{1}\int_{s}^{1}(s+x(t-s))^{n-1}\left(\int_{0}^{s}(s_{i}+x(1-s_{i}))\,ds_{i}\right)^{n-1}(s+x(1-s))\,dtds.$
Here s+x(ts)s+x(t-s) is the contribution of a vertex $w\in A$ that is not $v_{A}$. If $x_{w}=s_{i}<s$ for some $w\in B$ that is not $v_{B}$, then the attached leaf in $C$ has contribution $s_{i}+x(1-s_{i})$. (Note that attached leaf vertex $u$ can have a value bigger than $t$ as $t$ was only a maximum among the vertices of $A$.) The last term s+x(1s)s+x(1-s) is simply the contribution of the leaf attached to $v_{B}$.
@@
(3.2) $I_{2}:=n^{2}\int_{0}^{1}\int_{s}^{1}t^{n-1}\left(\int_{0}^{t}(s_{i}+x(1-s_{i}))\,ds_{i}+\int_{t}^{s}x(1-s_{i})\,ds_{i}\right)^{n-1}x(1-s)\,dtds.$
Note that we can assume that none of the vertices of $B$ are active since the contribution of such a term would be 0. Here, the contribution of a vertex $w\in A\setminus\{v_{A}\}$ is simply $t$ as it can never be active since $t<s$. If we have a vertex $w\in B$ that is not $v_{B}$ with $x_{w}=s_{i}<s$, then we need to distinguish two cases. If $s_{i}<t$, then $w$ is not active, and so the attached leaf has a contribution $s_{i}+x(1-s_{i})$. If $s_{i}>t$ then, in order to make $w$ inactive the attached leaf $u$ should have a value $x_{u}>s_{i}$, but then it is automatically active and its contribution is $x(1-s_{i})$. The contribution of the leaf attached to $v_{B}$ is x(1s)x(1-s) as it has to be active to make $v_{B}$ inactive.
@@
$\lim_{n\to\infty}I_{1}^{1/n}$ $=\max_{0\leqslant s\leqslant t\leqslant 1}(s+x(t-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
$=\max_{0\leqslant s\leqslant 1}(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
The function $(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$ has either a maximum at $\left(1-\frac{1}{\sqrt{3}}\right)\frac{x}{x-1}$ if this is less than $1$, or at $s=1$. In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}$. In the second case, its value is $\frac{1}{2}(x+1)$.
@@
$\lim_{n\to\infty}I_{2}^{1/n}=\max_{0\leqslant t\leqslant s\leqslant 1}t\left((1-x)\frac{t^{2}}{2}+xt+x(s-t)-x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
Observe that at $s=t=1$ its value is $\frac{1}{2}(x+1)$. On the other hand, this is the maximum as dropping the condition $t\leqslant s$ would lead to the following chain of inequalities
$\lim_{n\to\infty}I_{2}^{1/n}$ $=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
$\leqslant\max_{0\leqslant t,s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
$=\max_{0\leqslant s\leqslant 1}\left(x\left(s-\frac{s^{2}}{2}\right)+\frac{1}{2}\right)$
$=\frac{1}{2}(x+1).$
So $\lim_{n\to\infty}I_{2}^{1/n}=\frac{1}{2}(x+1)$ which is always at most $\lim_{n\to\infty}I_{1}^{1/n}$. Then we have
@@
This proves the first statement.
The statement concerning $\widetilde{T}_{H_{n,n,n}}(0,x)$ is actually simpler. As before, let
$t=\max_{v\in A}x_{v}\ \ \ \text{and}\ \ \ s=\max_{v\in B}x_{v},$
and let $v_{A}$ and $v_{B}$ be the vertices, where these maximums are achieved. Observe that we only have to consider the case $t\leqslant s$, otherwise the vertex $v_{A}$ is active and $X_{v_{A}}=0$. If $t\leqslant s$, then for any $v\in A$ different from $v_{A}$ the value $x_{v}$ can be anything between $[0,t]$ without $v$ being active. If for a vertex $v\in B$ different from $v_{B}$ we have $x_{v}=s_{i}$, then its unique neighbor in $C$ cannot take bigger value than $s_{i}$, otherwise it becomes active and the contribution of this permutation would be 0. If $s_{i}<t$, then $v$ is not active, while if $s_{i}\in(s,t)$, then $v$ is active. Putting all these together, we get that
$\widetilde{T}_{H_{n,n,n}}(0,x)$ $=n^{2}\int_{0}^{1}\int_{0}^{s}t^{n-1}\left(\int_{0}^{t}s_{i}\,ds_{i}+x\int_{t}^{s}s_{i}\,ds_{i}\right)^{n-1}s\,dtds$
$=n^{2}\int_{0}^{1}\int_{0}^{s}t^{n-1}\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)^{n-1}s\,dtds.$
@@
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}$ $=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
$=\max_{0\leqslant t\leqslant 1}t\left(\frac{t^{2}}{2}+x\left(\frac{1}{2}-\frac{t^{2}}{2}\right)\right)$
This expression has a maximum either at $t=\left(\frac{x}{3(x-1)}\right)^{1/2}$ if $x\geqslant\frac{3}{2}$ or at $1$ if $x\leqslant\frac{3}{2}$. In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}$. In the second case, its value is $\frac{1}{2}$. Since $\widetilde{T}_{H}(1,1)=1$ for any bipartite graph $H$ by definition, this completes the sketch of the proof. ∎
@@
There are two takeaways from this section. First, sometimes it is possible to compute the exponential growth of the permutation Tutte polynomial without explicitly computing it. Second, it is often possible to understand those permutations that contribute the most weight to the permutation Tutte polynomial. Concentrating on these permutations might give a better lower or upper bound on the permutation Tutte polynomial.